# Support Graph Preconditioners for Off-Lattice Cell-Based Models

### v1

@@ 3.1 Vaidya’s Preconditioners
Vaidya’s second class of preconditioners builds on the first by adding edges back to an MST. Given a parameter $t$, we split $T$ into $t$ disjoint subtrees of roughly the same size where each subtree has at most $m/t$ vertices. Then we add the maximum weight edge in $\mathfrak{C}$ between each pair of subtrees if they are connected in $\mathfrak{C}$ but not in $T$. Let $\mathfrak{T}^{\prime}$ be the augmented support graph and $P^{\prime}$ its block Laplacian. We call $P^{\prime}$ an augmented MST preconditioner of $\Gamma$. The theoretically optimal value of $t$ is approximately $n^{1/4}$ [7].
The only step left to define is how we generate the support graph. Prim’s algorithm is the best choice for finding an MST because it tells us, at no extra cost, how to permute the rows and columns of the block Laplacian to generate zero fill during factorization. We prove this in the next subsection. The time complexity of Prim’s algorithm is $\mathcal{O}(m\log n)$. Augmenting an MST is straightforward once it has been decomposed into subtrees. A simple partitioning algorithm is presented in [7, TreePartition] and a more sophisticated one in [35].
In practice, it is often beneficial to give the preconditioner the same diagonal blocks as the original matrix. This has the intuitive benefit of capturing all the information contained in a Jacobi preconditioner. Doing so invalidates our analysis in the next section of the minimum eigenvalue of the preconditioned system, but it also decreases the maximum eigenvalue. While this choice is theoretically unfounded, it can be helpful in the reality of finite precision arithmetic (see Section 5).
3.2 The Elimination Game
@@
In the game, all the vertices of a graph are eliminated in some order (e.g., see Fig. 2 where the elimination order is according to the vertex labels). When a vertex $v$ is eliminated, fill edges are constructed so that every uneliminated neighbor of $v$ (connected by either an original or fill edge) becomes pairwise adjacent. In other words, the uneliminated neighbors of $v$ become a clique. The set of fill edges is in bijection with the set of fill entries that would be created during Gaussian elimination or decomposition. Note that the elimination game is never explicitly implemented but rather implicitly performed. Using this game, we can easily analyze the amount of fill created by a given permutation. The general problem of minimizing fill is NP-hard [40], but trees can easily be ordered to produce no fill.
Figure 2: An example of the elimination game being played on the same graph with different orderings. Graph edges and matrix entries are in solid black and denoted by crosses respectively. Fill edges and entries are in dotted blue and blue circles respectively.
Lemma 3.1
@@
Proof: The first vertex eliminated must be a leaf. This produces no fill and yields another rooted tree. Inductively, the hypothesis demands that each vertex must be a leaf when it is eliminated, otherwise, it would have children to eliminate first. Since eliminating leaves produces no fill, we get the desired result. $\blacksquare$
Corollary 3.2
@@
3.3 Solving Preconditioners
Matrix inverses are almost never explicitly computed. Instead, large matrices are decomposed into the product of diagonal and triangular matrices in which solving systems is easy. For a symmetric positive definite matrix $M$, the Cholesky decomposition finds a lower triangular matrix $L$ such that $M=LL^{T}$. We instead opt for the similar $LDL^{T}$ decomposition in which $D$ is a diagonal matrix and $L$ has only ones on its diagonal. This is preferable for block matrices because it avoids computing matrix square roots. This section provides a general decomposition algorithm for symmetric positive definite block matrices and adapts it to the special case of nonsingular block Laplacians of trees.
@@ 3.3 Solving Preconditioners
1:function LDLT($A$: $n\times n$ symmetric positive definite block matrix)
@@
1:function TreeLDLT($A$: nonsingular block Laplacian of a rooted tree)
@@
Proof: We assume the validity of Algorithm 1 and prove its equivalence to Algorithm 2 in this special case. Say that the rows and columns of $A$ are ordered as specified on line 3 of Algorithm 2, so $i<j$ means that vertex $i$ is at least as far from the root as vertex $j$. In this case, we say that $i$ is younger than $j$ and that $j$ is older than $i$. Note that vertices $i$ and $j$ being connected is equivalent to $\uline{A_{ij}}$ being nonzero. Lines 5–7 in both algorithms behave identically because the only vertices younger than $i$ that are also connected to $i$ are its children.
We want to show the equivalence of lines 10–16 in Algorithm 1 to lines 10–11 in Algorithm 2. We prove by induction on $i$ that, in Algorithm 1, $\uline{L_{ji}}$ is nonzero only if $i$ is connected to $j$ for all $j>i$. The hypothesis holds for the first vertex because $Y$ is guaranteed to be zero as there are no younger vertices. Now assume that the hypothesis holds for all vertices younger than some $i$. If $Y$ is nonzero, then there exists an older vertex $j$ and a younger vertex $k$ such that $k$ is connected to both $i$ and $j$. However, the only vertex older than $k$ that it is connected to is its unique parent. This implies that $i=j$, which contradicts line 10. Therefore, $Y$ is zero and the hypothesis is true, meaning the only $j$ we need to consider is the parent of $i$. $\blacksquare$
When there is zero fill, no additional memory allocation is required for $L$ even when using a matrix-free or sparse matrix implementation. However, this is not the case for augmented MST preconditioners. Algorithm 2 is no longer valid and extra fill entries need to be stored. Another benefit of the standard MST preconditioner is that solving systems in the decomposed block Laplacian takes linear time. We provide algorithms for this in Section B.1.
This begs the question of whether there exists an effective augmentation strategy that does not create fill. This restricts us to only add back edges between siblings in the MST. However, we suspect that this would not yield promising results for reasons we formalize in the next section. Intuitively, we want to add back edges that drastically reduce the distance between pairs of vertices, but sibling edges do a poor job of this.
Finally, it is known that block LU factorization for symmetric positive definite matrices is stable as long as the matrix is well-conditioned [9, 19].
4 Block-Structured Support Graph Theory
@@
We use $\succeq$ to represent the Loewner order. For matrices $A$ and $B$, we say $A\succeq B$ if and only if $A-B$ is positive semidefinite. A fact that we use without proof is that the Loewner order is a partial order on symmetric matrices. We denote the set of finite generalized eigenvalues of a pair of matrices $(A,B)$ by $\lambda(A,B)$. That is to say, $\lambda(A,B)$ is the set of numbers $\lambda$ such that there exists a vector $\mathbf{x}$ where $A\mathbf{x}=\lambda B\mathbf{x}$. If $B$ is a preconditioner for $A$, then the condition number of the preconditioned system $B^{-1}A$ is precisely the ratio of the extremal finite generalized eigenvalues $\lambda_{\mathrm{max}}(A,B)/\lambda_{\mathrm{min}}(A,B)$.
@@
Proof: Observe that $L_{G}=L_{H}+L_{K}$ where $L_{K}$ is the block Laplacian of the graph $K=(V,E\setminus F,w-w^{\prime})$. Since $L_{G}-L_{H}=L_{K}$ is a block Laplacian, it is positive semidefinite. This implies that $\sigma(L_{H},L_{G})\leq 1$ and that $\lambda_{\mathrm{min}}(L_{G},L_{H})\geq 1$. $\blacksquare$
@@
Proof: Since $A$ is symmetric positive semidefinite, it is diagonalizable and we can write $A=PDP^{-1}$ where $D$ is the diagonal matrix whose entries are the eigenvalues of $A$. For concision, let $\lambda=\lambda_{\mathrm{max}}(A)$. Observe
@@
By the definition of $\lambda$, the middle factor of $\lambda I-D$ has all nonnegative entries, so the whole difference is positive semidefinite. Similar logic applies for the minimum eigenvalue. $\blacksquare$
@@
Proof: Let $C=k\hat{B}-A$ and define $\hat{C}=\mathrm{diag}(A^{-1})C$ which is a block-structured matrix in which the blocks are either nonzero or a multiple of the identity matrix. Since $\mathrm{diag}(A^{-1})$ is positive definite, it follows that $C$ is semi-positive definite if $\hat{C}$ is positive semidefinite. We prove that $\hat{C}$ is positive semidefinite using induction as in the non-block-structured case in [3, Lemma 2.7]. The only difference is that we use a block-structured symmetric Gaussian elimination, meaning the $i$-th elimination step is
@@
Since the matrix $\hat{C}$ has nonnegative values on its diagonal, this shows that $\hat{C}$ is positive semidefinite. $\blacksquare$
Lemma 4.8
@@
Proof: Let $\lambda=\lambda_{\mathrm{max}}(AB^{-1})$. When $A=B$, we have that $\lambda=1$ and the statement is equivalent to Lemma 4.7. When $A\neq B$, we can reduce to the case of equality by multiplying $\hat{B}$ by $\mathrm{diag}(AB^{-1})$. This yields
@@
Next we show that $k\cdot\lambda\hat{B}\succeq k\cdot\mathrm{diag}(AB^{-1})\hat{B}$. This is equivalent to showing that $\lambda I\succeq AB^{-1}$, which is true by Lemma 4.6. The rest follows from the transitivity of the Loewner order. $\blacksquare$
@@
Lemma 4.9 (congestion-dilation lemma)
@@
Proof: Let $D=\min_{i}\{\lambda_{\mathrm{min}}(B_{i})\}\cdot I$. Decompose
@@
The first summand is positive semidefinite by Lemma 4.8. The diagonal blocks of $\hat{B_{2}}$ are positive semidefinite and the nonzero off-diagonal blocks are negative semidefinite by Lemma 4.6. This and the fact that the block row sums are zero means the second summand is a block Laplacian, so it is positive semidefinite. $\blacksquare$
Under the support of the path represented by $\hat{B}$, we call $\lambda$ the congestion of the edge represented by $\hat{A}$, and $k$ is its dilation. A more concise statement of the lemma is that $\sigma(\hat{A},\hat{B})$ is bounded above by the product of the congestion and dilation. We use the this lemma to prove an upper bound on the maximum eigenvalue of a block Laplacian with an MST preconditioner. The following proofs make use of the fact that
@@
Theorem 4.10
Let $G=(V,E,w)$ be a matrix-weighted graph with MST $T$ weighted by minimum eigenvalues. Let $L_{G}$ and $L_{T}$ be their block Laplacians and let $\kappa$ be the maximum condition number of all the edge weights in $G$. Then
@@
Proof: For every edge $e=(u,v)$, let $p(e)$ be the path in $T$ from $u$ to $v$ that uses at least $1/m$ fraction of each edge weight. We can write
@@
The maximum possible length (edge count) of each $p(e)$ is $n-1$. This is the dilation. Since $T$ is an MST, the minimum eigenvalue of each edge weight in $p(e)$ is at least that of $w(e)$. This means that the congestion is at most $\kappa m$. The congestion-dilation lemma yields the desired result. $\blacksquare$
@@
Proof: We perform a similar decomposition to the previous proof. Let $S_{1},\ldots,S_{t}$ be the subtrees of $T^{\prime}$. For each edge $e=(u,v)$, a path from $u$ to $v$ in $T^{\prime}$ is not necessarily unique. If both $u$ and $v$ are in the same $S_{i}$, let $p(e)$ be the unique path from $u$ to $v$ contained in $S_{i}$. If $u$ is in $S_{i}$ and $v$ is in $S_{j}$ with $i\neq j$, let $p(e)$ be the concatenation of the following paths: the unique path in $S_{i}$ from $u$ to the endpoint of the edge that connects $S_{i}$ and $S_{j}$, that edge itself, and the unique path in $S_{j}$ from the other endpoint to $v$.
@@
In the worst case, one of these paths may go across all $dn/t-1$ edges in one subtree, the edge connecting it to another subtree, and all $dn/t-1$ edges in the other subtree. Therefore, the dilation is less than $2(dn/t)$ and the rest follows from the congestion-dilation. $\blacksquare$
@@
5 Numerical Benchmarks
Figure 3: The average conjugate gradient convergence of $25$ experiments, each of $N=10000$ randomly placed cells in a large 3D-sphere. The resulting block-structured linear systems are solved using the conjugate gradient method with five different preconditioning strategies: (1) the identity; (2) Jacobi precondtioner; (3) block Jacobi precondtioner; (4) the support-graph preconditioner; and (5) the row-support graph preconditioner. In the top row $\gamma_{\mathrm{med}}=3\times 10^{5}$, $\gamma_{\parallel}=2\times 10^{6}$ and $\gamma_{\perp}=8\times 10^{6}$. In the middle row $\gamma_{\mathrm{med}}=3\times 10^{4}$, $\gamma_{\parallel}=2\times 10^{6}$ and $\gamma_{\perp}=8\times 10^{6}$. In the bottom row $\gamma_{\mathrm{med}}=3\times 10^{4}$, $\gamma_{\parallel}=2\times 10^{7}$ and $\gamma_{\perp}=8\times 10^{7}$.

### v2

@@ 3.1. Vaidya’s Preconditioners
Vaidya’s second class of preconditioners builds on the first by adding edges back to an MST. Given a parameter $t$, we split an MST into $t$ disjoint subtrees of roughly the same size where each subtree has at most $m/t$ vertices. Then we add the maximum weight edge in $\mathfrak{C}$ between each pair of subtrees if they are connected in $\mathfrak{C}$. Let $\mathfrak{T}^{\prime}$ be an augmented MST and $P^{\prime}$ its block Laplacian. We call $P^{\prime}$ an augmented MST preconditioner of $\Gamma$. The theoretically optimal value of $t$ is approximately $n^{1/4}$ [6].
The only step left to define is how we generate the support graph. Prim’s algorithm is the best choice for finding an MST because it tells us, at no extra cost, how to permute the rows and columns of the block Laplacian to generate zero fill during factorization. We prove this in the next subsection. The time complexity of Prim’s algorithm is $\mathcal{O}(m\log n)$. Augmenting an MST is straightforward once it has been decomposed into subtrees. A simple partitioning algorithm is presented in [6, TreePartition] and a more sophisticated one in [38].
3.2. The Elimination Game
@@
In the game, all the vertices of a graph are eliminated in some order (e.g., see Figure 2 where the elimination order is according to the vertex labels). When a vertex $v$ is eliminated, fill edges are constructed so that the uneliminated neighbors of $v$ (connected by either an original or fill edge) become pairwise adjacent. In other words, the uneliminated neighbors of $v$ become a clique. The set of fill edges is in bijection with the set of fill entries that would be created during Gaussian elimination or decomposition. Note that the elimination game is never explicitly implemented but rather implicitly performed. Using this game, we can easily analyze the amount of fill created by a given permutation. The general problem of minimizing fill is NP-hard [43], but trees can easily be ordered to produce no fill.
Figure 2. An example of the elimination game being played on the same graph with different orderings. Graph edges and matrix entries are in solid black and denoted by crosses respectively. Fill edges and entries are in dotted blue and blue circles respectively.
Lemma 3.1.
@@
Proof.
The first vertex eliminated must be a leaf. This produces no fill and yields another rooted tree. Inductively, the hypothesis demands that each vertex must be a leaf when it is eliminated, otherwise, it would have children to eliminate first. Since eliminating leaves produces no fill, we get the desired result. ∎
Corollary 3.2.
@@
3.3. Solving Preconditioners
Matrix inverses are seldom computed explicitly. Instead, large matrices are decomposed into the product of diagonal and triangular matrices in which solving systems is easy. For a symmetric positive definite matrix $M$, the Cholesky decomposition finds a lower triangular matrix $L$ such that $M=LL^{T}$. We instead opt for the similar $LDL^{T}$ decomposition in which $D$ is a diagonal matrix and $L$ has only ones on its diagonal. This is preferable for block matrices because it avoids computing matrix square roots. This section provides a general decomposition algorithm for symmetric positive definite block matrices and adapts it to the special case of nonsingular block Laplacians of trees.
@@ 3.3. Solving Preconditioners
1:function LDLT(AA: $n\times n$ symmetric positive definite block matrix)
@@
1:function TreeLDLT(AA: nonsingular block Laplacian of a rooted tree)
@@
Proof.
We assume the validity of Algorithm 1 and prove its equivalence to Algorithm 2 in this special case. Say that the rows and columns of $A$ are ordered as specified on line 3 of Algorithm 2, so $i<j$ means that vertex $i$ is at least as far from the root as vertex $j$. In this case, we say that $i$ is younger than $j$ and that $j$ is older than $i$. Note that vertices $i$ and $j$ being connected is equivalent to $\uline{A_{ij}}$ being nonzero. Lines 5–7 in both algorithms behave identically because the only vertices younger than $i$ that are also connected to $i$ are its children.
We want to show the equivalence of lines 10–16 in Algorithm 1 to lines 10–11 in Algorithm 2. We prove by induction on $i$ that, in Algorithm 1, $\uline{L_{ji}}$ is nonzero only if $i$ is connected to $j$ for all $j>i$. The hypothesis holds for the first vertex because $Y$ is guaranteed to be zero as there are no younger vertices. Now assume that the hypothesis holds for all vertices younger than some $i$. If $Y$ is nonzero, then there exists an older vertex $j$ and a younger vertex $k$ such that $k$ is connected to both $i$ and $j$. However, the only vertex older than $k$ that it is connected to is its unique parent. This implies that $i=j$, which contradicts line 10. Therefore, $Y$ is zero and the hypothesis is true, meaning the only $j$ we need to consider is the parent of $i$. ∎
When there is zero fill, no additional memory allocation is required for $L$ even when using a matrix-free or sparse matrix implementation. However, this is not the case for augmented MST preconditioners. Algorithm 2 is no longer valid and extra fill entries need to be stored for a complete factorization. Other options for solving systems in the augmented MST preconditioner include using an incomplete factorization with a fill-reducing algorithm like one of those mentioned in Section 3.2, or making a nested call to the conjugate gradient method with the standard MST as the preconditioner.
This begs the question of whether there exists an effective augmentation strategy with which complete factorizations do not create fill. This restricts us to only add back edges between siblings in the MST. However, we suspect that this would not yield promising results for reasons we formalize in the next section. Intuitively, we want to add back edges that drastically reduce the distance between pairs of vertices (i.e., reduce the stretch), but sibling edges do a poor job of this.
Another benefit of the standard MST preconditioner is that solving systems in the decomposed block Laplacian takes linear time. We provide algorithms for this in Section B.1. Finally, it is known that block $LU$ factorization for symmetric positive definite matrices is stable as long as the matrix is well-conditioned [8, 19].
4. Block-Structured Support Graph Theory
@@
We use $\succeq$ to represent the Loewner order. For matrices $A$ and $B$, we say $A\succeq B$ if and only if ABA-B is positive semidefinite. A fact that we use without proof is that the Loewner order is a partial order on symmetric matrices. We denote the set of finite generalized eigenvalues of a pair of matrices $(A,B)$ by $\lambda(A,B)$. That is to say, $\lambda(A,B)$ is the set of numbers $\lambda$ such that there exists a vector $\mathbf{x}$ where $A\mathbf{x}=\lambda B\mathbf{x}$. If $B$ is a preconditioner for $A$, then the condition number of the preconditioned system $B^{-1}A$ is precisely the ratio of the extremal finite generalized eigenvalues $\lambda_{\mathrm{max}}(A,B)/\lambda_{\mathrm{min}}(A,B)$.
@@
Proof.
Observe that $L_{G}=L_{H}+L_{K}$ where $L_{K}$ is the block Laplacian of the graph $K=(V,E\setminus F,w-w^{\prime})$. Since $L_{G}-L_{H}=L_{K}$ is a block Laplacian, it is positive semidefinite. This implies that $\sigma(L_{H},L_{G})\leq 1$ and that $\lambda_{\mathrm{min}}(L_{G},L_{H})\geq 1$. ∎
@@
Proof.
Since $A$ is symmetric positive semidefinite, it is diagonalizable and we can write $A=PDP^{-1}$ where $D$ is the diagonal matrix whose entries are the eigenvalues of $A$. For concision, let $\lambda=\lambda_{\mathrm{max}}(A)$. Observe
@@
By the definition of $\lambda$, the middle factor of $\lambda I-D$ has all nonnegative entries, so the whole difference is positive semidefinite. Similar logic applies for the minimum eigenvalue. ∎
@@
Proof.
Let $C=k\hat{B}-A$ and define $\hat{C}=\mathrm{diag}(A^{-1})C$ which is a block-structured matrix in which the blocks are either nonzero or a multiple of the identity matrix. Since $\mathrm{diag}(A^{-1})$ is positive definite, it follows that $C$ is semi-positive definite if $\hat{C}$ is positive semidefinite. We prove that $\hat{C}$ is positive semidefinite using induction as in the non-block-structured case in [2, Lemma 2.7]. The only difference is that we use a block-structured symmetric Gaussian elimination, meaning the $i$-th elimination step is
@@
Since the matrix $\hat{C}$ has nonnegative values on its diagonal, this shows that $\hat{C}$ is positive semidefinite. ∎
Lemma 4.8.
@@
Proof.
Let $\lambda=\lambda_{\mathrm{max}}(AB^{-1})$. When $A=B$, we have that $\lambda=1$ and the statement is equivalent to Lemma 4.7. When $A\neq B$, we can reduce to the case of equality by multiplying $\hat{B}$ by $\mathrm{diag}(AB^{-1})$. This yields
@@
Next we show that $k\cdot\lambda\hat{B}\succeq k\cdot\mathrm{diag}(AB^{-1})\hat{B}$. This is equivalent to showing that $\lambda I\succeq AB^{-1}$, which is true by Lemma 4.6. The rest follows from the transitivity of the Loewner order. ∎
@@
Lemma 4.9 (congestion-dilation lemma).
@@
Proof.
Let $D=\min_{i}\{\lambda_{\mathrm{min}}(B_{i})\}\cdot I$. Decompose
@@
The first summand is positive semidefinite by Lemma 4.8. The diagonal blocks of $\hat{B_{2}}$ are positive semidefinite and the nonzero off-diagonal blocks are negative semidefinite by Lemma 4.6. This and the fact that the block row sums are zero mean the second summand is a block Laplacian, so it is positive semidefinite. ∎
[…]
@@
Theorem 4.10.
Let $G=(V,E,w)$ be a matrix-weighted graph and $T$ its MST weighted by minimum eigenvalues. Let $L_{G}$ and $L_{T}$ be their block Laplacians and let $\kappa$ be the maximum condition number of all the edge weights in $G$. Then
@@
Proof.
For every edge $e=(u,v)$, let $p(e)$ be the path in $T$ from $u$ to $v$ that uses at least $1/m$ fraction of each edge weight. We can write
@@
The maximum possible length (edge count) of each $p(e)$ is n1n-1. This is the dilation. Since $T$ is an MST, the minimum eigenvalue of each edge weight in $p(e)$ is at least that of $w(e)$. This means that the congestion is at most $\kappa m$. The congestion-dilation lemma yields the desired result. ∎
@@
Proof.
We perform a similar decomposition to the previous proof. Let $S_{1},\ldots,S_{t}$ be the subtrees of $T^{\prime}$. For each edge $e=(u,v)$, a path from $u$ to $v$ in $T^{\prime}$ is not necessarily unique. If both $u$ and $v$ are in the same $S_{i}$, let $p(e)$ be the unique path from $u$ to $v$ contained in $S_{i}$. If $u$ is in $S_{i}$ and $v$ is in $S_{j}$ with $i\neq j$, let $p(e)$ be the concatenation of the following paths: the unique path in $S_{i}$ from $u$ to the endpoint of the edge that connects $S_{i}$ and $S_{j}$, that edge itself, and the unique path in $S_{j}$ from the other endpoint to $v$.
@@
In the worst case, one of these paths may go across all dn/t1dn/t-1 edges in one subtree, the edge connecting it to another subtree, and all dn/t1dn/t-1 edges in the other subtree. Therefore, the dilation is less than $2(dn/t)$ and the rest follows from the congestion-dilation lemma. ∎
@@
5. Numerical Benchmarks
Figure 3. Preconditioner convergence behavior for different cell arrangements: hexagonal lattices with varying disorder (top row), random packing (bottom left), and bridged spheroids (bottom right). The insets show cross-sections through the 3D cell configurations. The red dashed line in each plot indicates the solver relative tolerance $10^{-5}$. Legend: Identity/no preconditioner (orange dotted line); block Jacobi (green long dash-dot); diagonally-shifted block IC(0) (dashed red); block Gauss-Seidel (cyan dash-dot); MST preconditioner (solid blue); augmented MST (gray circular markers). Top Row: Cells arranged on a three-dimensional hexagonal lattice with positional noise of mean zero and standard deviation $0.3r$ applied to each cell. Each configuration contains $n\approx 50\,000$ cells. The edge-to-vertex ratio is reported for each configuration. Left: $\gamma_{\mathrm{med}}=3\times 10^{4}$. Right: $\gamma_{\mathrm{med}}=3\times 10^{3}$. $\kappa$ in the title is the full friction matrix condition number. Bottom Left: n=50 000n=50\,000 cells randomly packed in a spheroidal domain with $\gamma_{\mathrm{med}}=3\times 10^{4}$. Bottom Right: Two densely packed spheroids connected by a bridge of randomly placed cells, $n\approx 50\,000$ total, $\gamma_{\mathrm{med}}=3\times 10^{4}$.
Figure 4. Performance comparison of preconditioners for block-structured linear systems arising from $n$ cells arranged in a hexagonal lattice with positional noise (30% of cell radius). Each data point represents the average of 5 independent experiments. The friction parameters are $\gamma_{\mathrm{med}}=3\times 10^{4}$ with $\gamma_{\parallel}=2\times 10^{6}$ and $\gamma_{\perp}=8\times 10^{7}$. (Left) Total wall-clock time as a function of problem size $n$ for solving the linear system to relative tolerance $10^{-5}$. Times include all computational costs: matrix assembly (where required), preconditioner construction and factorization, and conjugate gradient iterations. The MST preconditioner achieves a speedup of $3.5\times$ compared to unpreconditioned CG at $n=10^{5}$. The direct solver (Eigen LDLT) excels for $n<10^{3}$ but becomes catastrophically slow for larger problems. Preconditioners tested: (1) identity (no preconditioning), (2) block Jacobi with $LDL^{T}$ factorization of diagonal blocks, (3) MST with tree construction and factorization, (4) augmented-MST with tree construction, augmentation and nested (flexible) conjugate gradient (inner relative tolerance $\delta=0.1$), (5) IC(0) with diagonal shift and incomplete factorization, (6) block Gauss-Seidel with setup, and (7) Eigen’s sparse $LDL^{T}$ direct solver including assembly and factorization. The MST preconditioner achieves the fastest total solution times. (Right) Iteration counts for the same preconditioners as a function of $n$.